OpenStudy (anonymous):

sinx=-2sinxcosx help mee? :D

OpenStudy (anonymous):

i know im not suppose to divide by sinx but idk how to do it otherwise

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

@TillyAsks

OpenStudy (anonymous):

oh u already here :P

OpenStudy (anonymous):

Okay factor out sinx

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

like sinx=sinx(-2cosx)

OpenStudy (anonymous):

1+2cos x = 0

OpenStudy (anonymous):

u divided by sinx? u cant do that though because sinx can be equal to 0

OpenStudy (anonymous):

Gah this is complicated one see let me pen and paper it ..can use my brain

OpenStudy (anonymous):

yea i did what u just did and got points off cuz aprently doing it this way will only get u half answers because sinx can be 0 :(

OpenStudy (raden):

sinx = -2sinxcosx sinx + 2sinxcosx = 0 sinx(1 + 2cosx) = 0 the zeroes is when each factor equals zero sinx = 0 solve for x and 1 + 2cosx = 0 solve for x

OpenStudy (anonymous):

thank u i love u!

OpenStudy (raden):

really, oh my good :* haha

OpenStudy (anonymous):

Ha Sorry :c